Expand and factor the expression x(x−1)−30.
(b) Solve the equation x(x−1)=30. If there is more than one correct answer, enter your answers as a comma separated list. If there are no answers, enter NONE.

Answer by DrBeeee(684)

Given:
(1) x*(x-1)-30 = 0 which expands to
(2) x^2 -x -30 = 0 which factors into
(3) (x-6)*(x+5) = 0
The roots of (3) are the solution to the equation
(4) x = (-5,6)
Let's check the solution using (1).
Is (-5*(-5-1)-30 = 0)?
Is (-5*(-6)-30 = 0)?
Is (30 - 30 = 0)?
Is (0 = 0)? Yes
Is (6*(6-1)-30 = 0)?
Is (6*5 - 30 = 0)?
Is (30 - 30 = 0)?
Is (0 = 0)? Yes
Answer: The solution set is {-5,6}